OpenStudy (anonymous):

The following 3 items relate to a 50,000-watt radio station that transmits 750 kHz(WSB). 1. Determine the number of joules transmitted per second. A. 750 B. 50 000 C. 750, 000 D. 1.6x10^19 E. 8x10^23 3. calculate the energy of a single photon at the transmitted frequency. A. 5.0x10^-31 j B. 1.4x10^-29 j C. 5.0x10^9-28 j D. 1.4x10^-19 j E. 7.5x10^-5 j 2. what is the wavelength of the radio waves emitted from this radio station in meters per cycle.( hint : use the wave equation) A. 1.4x10^-29 B. 5 C. 10.5 D. 75 E. 400

OpenStudy (souvik):

1 watt power=1 joule work per second

OpenStudy (souvik):

the energy of a photon E=hf where h=plank's constant and f=frequency

OpenStudy (anonymous):

I don't get it

OpenStudy (souvik):

first one?

OpenStudy (anonymous):

yes

OpenStudy (souvik):

if i do 1 joule work in one second...then you can say my power is 1 Watt...

OpenStudy (souvik):

similarly the power of radio station is 50,000 Watt that means it transmitted 50,000 joules per second

OpenStudy (anonymous):

okay thank you , number 2 please ?

OpenStudy (souvik):

E=hf

OpenStudy (souvik):

the energy of a photon E=hf where h=plank's constant and f=frequency

OpenStudy (theeric):

@smartG Are you on the second one - number 3?

OpenStudy (theeric):

"calculate the energy of a single photon at the transmitted frequency." I'll quote souvik, as his last comments answer that. These comments: "\(E=hf\)" and "the energy of a photon \(E=hf\) where \(h\)=plank's constant and \(f\)=frequency" Planck's constant is \(6.62606957 × 10^{-34}\ m^2\ kg / s\).

OpenStudy (anonymous):

so you just plug it in to he formula ?

OpenStudy (souvik):

yeah

OpenStudy (anonymous):

how do I know what to plug in for what ?

OpenStudy (souvik):

As Eric wrote h=6.6*10^-34 SI unit f=750,000 Hz

OpenStudy (anonymous):

i didn't get one of the answer choices

OpenStudy (souvik):

it is 4.9*10^-28 ~5*10^-28

OpenStudy (anonymous):

so # 2 is c?

OpenStudy (theeric):

I got \[E=hf\]\(6.636\times10^{-31}\ \ \times\ \ 750000\ \ =\ \ 4.9695 \times 10^{-28}\) http://www.wolframalpha.com/input/?i=6.626e-34+*+750e3 If you round this to two significant figures, you find your answer.

OpenStudy (theeric):

Might have been a math error.

OpenStudy (theeric):

As for the next problem... The wave equation it speaks of is \(v=f\lambda\), where \(v\) is the velocity at which the wave propagates, \(f\) is the frequency, and \(\lambda\) is the wavelength you seek. So, you can rearrange that, solving for \(\lambda\). Do you know the velocity of a electromagnetic wave? (Same for all electromagnetic waves, like radio waves, light, ultraviolet radiation, and other names for other frequencies). You'll get it! :)